OpenStudy (anonymous):

x+y=0 x-y=k If the ordered pair (1, p), where x=1 and y=p, is a solution of both equations above, then what is the value of k?

OpenStudy (anonymous):

saki did u try to plug it in?

OpenStudy (anonymous):

since x=1 right? plug it in to x+y=0 which will be like 1+y=0 and ull know what y is

OpenStudy (anonymous):

Yup! I plugged it in

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Wait, I get two. Because for the top equation 1+p=0, I moved the 1 to the other side, so it became p=-1 and then I plugged it into 1-p=k. So it became 1-(-1)=2

OpenStudy (anonymous):

its 1+y=0 so u move the 1 to other side which will be -1 when he moves to other side so it'll be like